OpenStudy (anonymous):

Simplify the expression. State any restrictions on the variable. 1-x/x^2+5x-6

OpenStudy (anonymous):

\[1-x \over x^2+5x-6\]

OpenStudy (anonymous):

x^2 + 5x -6 =(x+6)(x-1) (1-x)/((x^2+5x-6)) = -(x-1)/((x+6)(x-1)) = -1/(x+6) (x+6)=0 x=-6 denominator cannot be zero otherwise, the fraction is undefined. so x cannot be -6